# Sa2VA: Marrying SAM2 with LLaVA for Dense Grounded Understanding of Images and Videos

### v1

@@ 2 Related Work
Referring Segmentation. This task aims to output specific masks (for image inputs) or tracked masks (for video input) driven by language description. Earlier works [91, 52, 38, 13, 80, 101] explore various fusion modules to achieve better performance. Then, several stronger models [84] adopt transformer-based methods to achieve unified segmentation and tracking in video. Equipped with LLMs, several recent works [40, 99, 83, 65] propose more complex referring tasks, including reasoning referring or joint mask and caption generation. For example, LISA [40] involves reasoning-based segmentation. Then, GLaMM [66] annotates a new dataset and proposes region-level caption and segmentation tasks. Meanwhile, several recent works explore joint instruction tuning on referring segmentation and conversation. Our method extends these studies into the video domain with the help of SAM-2, while keeping stronger performance on both image/video referring tasks and VQA tasks.
Video Segmentation and Grounding. Current video segmentation methods [47, 34, 105] focus on segmenting and tracking pixels in closed sets. Several works [25, 103] explore open-vocabulary settings. However, the concepts are still limited compared with the knowledge space of LLMs. For video grounding, recent works [28] explore LLM into the joint understanding of video and audio. However, there is no fine-grained spatial-temporal modeling or enough image data for co-training, which makes the results unconvincing. Our model, Sa2VA, adds interactive chat and open-ended understanding to current dense video perception models and achieves strong VQA performance.
@@ 3 Method
Figure 2: Our proposed Sa2VA model. The model first encodes the input texts, visual prompts, images, and videos into token embeddings. These tokens are then processed through a large language model (LLM). The output text tokens are used to generate the [SEG] token and associated language outputs. The SAM-2 decoder receives the image and video features from the SAM-2 encoder, along with the [SEG] token, to generate corresponding image and video masks.
@@ 3.1 Unifying Multi-task Representations
Image/Video Chat and Grounded Caption Generation. For image and video chat tasks, given input text tokens $T_{i}$ and corresponding images $I_{i}$ or video $V_{i}$, the model outputs answer text $T_{o}$. For grounded caption generation tasks, the model jointly outputs the corresponding masks $M_{o}$ and aligned text $T_{o}$.
@@
Unified Task Representation. Existing works solve the above tasks via specific design models or partial unified models (different tasks have different model weights). In this work, we point out that all the above tasks can be unified as a one-shot instruction-tuning process, since we can leverage the flexibility of LLMs to handle various visual tokens. The overall process can be formulated as the following:
@@
For chat-only tasks, the model only outputs text tokens $T_{o}$. For referring segmentation tasks, the model outputs masks (image) or masklets (video) $M_{o}$. For grounded caption generation tasks, the model outputs both texts and masks. For different tasks, the input visual tokens are changing according to specific tasks. Since $M_{o}$ (whether image masks or masklets) can be controlled with SEG tokens, one SEG token represents one prompt input of SAM-2’s decoder. Thus, it is natural to put all these tasks into one shared encoder-and-decoder framework and co-train one LLM and fine-tune the SAM decoder.
Figure 3: Data annotation pipeline. Our proposed automatic data annotation pipeline consists of three stages: object/part-level, scene-level, and video-level text expression annotation. We use different colors in the final expression to highlight the information derived from each stage. Best view on screen and zoom out.
@@ 3.2 Sa2VA Framework
Pre-trained MLLMs. We adopt pre-trained LLaVA-like models as the MLLMs. It contains one visual encoder, one visual projection layer, and one LLM. The visual encoder takes input images, video, and sub-images as inputs. The visual projection layer maps inputs into visual tokens. These tokens, combined with the input text tokens, are the input of LLMs and perform output text token prediction. Note that we adopt pre-trained MLLMs following previous works [99, 66, 40, 85]. For both image and video chat datasets, we follow the same pipeline [9, 2] without further modification.
@@
Tuning SAM-2 Decoder via SEG Tokens. We connect SAM-2 and the MLLM via the special token "[SEG]". The hidden states of the "[SEG]" token are used as a new type of prompt and fed into SAM-2’s Decoder, where they are decoded into segmentation masks. The hidden states of "[SEG]" can be seen as a novel spatial-temporal prompt for SAM-2. SAM-2 segments the corresponding object mask in image and video based on the spatial-temporal prompt. During training, the SAM-2 decoder can be tuned to understand the spatial-temporal prompt, and gradients can be backpropagated through the "[SEG]" token to the MLLM, allowing it to output the spatial-temporal prompt better.
Utilizing SAM-2 knowledge for mask tracking. For Ref-VOS tasks, we design a simple framework to achieve strong results on public benchmarks. In particular, for giving input video, we adopt a "[SEG]" token to generate the masks of the key frames. Then, we use the memory encoded by the key frame features to generate the mask for the remaining frames. We detailed this process in the appendix.
@@ 3.3 Ref-SAV Dataset and Benchmark
    Object/part-level annotation. We first select the frame with the largest object area from the video and mask out the non-object pixels. The cropped and full images are then fed separately into InternVL2-76B [9] to generate detailed descriptions. The descriptions are processed by Qwen2-72B [86] for consistency checking, with conflicting descriptions discarded.
@@
    Scene-level annotation. Yellow contours are used to highlight the object in the image. Both the image and the object/part-level description from the previous stage are sent into InternVL2-76B [9] to generate a detailed object description that includes relationships to the scene and surrounding objects.
@@
Table 2: Experiment results on image/video referring segmentation benchmarks and image/video chat benchmarks.
@@
RefCOCO [36] RefCOCO+ [36] RefCOCOg [90] MeViS [13] Ref-DAVIS17 [38] ReVOS [85] MME [19] MMBench [59] SEED-Bench [41] Video-MME [20] MMBench-Video [17] GCG [66]
LLAVA-1.5-13B [54] - - - - - - 1531 68.8 70.1 - - -
Video-LLaVA-7B [51] - - - - - - - 60.9 - 39.9 1.03 -
LLaMA-VID-7B [50] - - - - - - 1521 65.1 59.9 - 1.08 -
mPLUG-Owl3-8B [89] - - - - - - - 77.6 - 53.5 1.35 -
InternVL2-8B [9] - - - - - - - 81.7 76.2 54.0 1.28 -
PixelLM-7B [68] 73.0 66.3 69.3 - - - 309/135 17.4 - - - -
LaSagnA [77] 76.8 66.4 70.6 - - - 0/0 0.0 - - - -
LISA-7B [40] 74.1 62.4 66.4 - - - 1/1 0.4 - - - -
GLaMM-7B [66] 79.5 72.6 74.2 - - - 14/9 36.8 - - - 28.9
LLaVA-G-7B [96] 77.1 68.8 71.5 - - - - - - - - -
GSVA-13B [83] 79.2 70.3 75.7 - - - - - - - - -
OMG-LLaVA-7B [99] 78.0 69.1 72.9 - - - 1177/235 47.9 56.5 - - 29.9
VISA-13B [85] 72.4 59.8 65.5 44.5 70.4 50.9 - - - - - -
Sa2VA-1B (Ours) 77.4 69.9 72.3 50.8 72.3 47.6 1381/405 68.3 64.8 39.9 1.07 23.8
Sa2VA-4B (Ours) 78.9 71.7 74.1 52.1 73.8 53.2 1536/530 77.3 73.3 50.4 1.23 28.2
Sa2VA-8B (Ours) 81.6 76.2 78.7 57.0 75.2 57.6 1617/511 81.6 75.1 52.1 1.34 31.0
Sa2VA-26B (Ours) 82.5 78.8 79.7 57.3 77.0 58.4 1691/538 83.7 76.8 52.6 1.45 33.5
@@ 4 Experiments
Baseline. We construct the baseline by combining the SOTA MLLM models like InternVL2 [9] and SAM2 [67]. Similar to previous works  [40, 66, 99], the segmentation mask is obtained by decoding the hidden state of the [SEG] token through SAM2’s decoder. Inspired by Mask2Former-VIS [10], an object shares the same [SEG] token throughout the video frames, allowing our model to handle both image and video referring segmentation tasks in a unified way. In addition, we further use more advanced MLLM models, InternVL2-5 to show the model scaling effect of Sa2VA.
@@
Table 4: Performance on image-level benchmarks. Our method achieves the best accuracy trade-off between image chat and referring segmentation datasets.
@@ 4.2 Ablation Studies
MLLM RefCOCO [36] RefCOCO+ [36] RefCOCOg [90] MeViS [13] Ref-DAVIS17 [38] MME [19] MMBench [59] SEED-Bench [41] AI2D [37] MMStar [7] SQAtest [60]
InternVL2.0-4B 80.4 74.3 76.7 52.1 73.8 1553/540 76.8 72.6 79.9 53.7 95.8
InternVL2.0-8B 81.9 76.5 78.9 57.0 75.2 1651/578 82.4 75.5 82.1 60.3 96.8
InternVL2.5-4B 82.4 77.6 79.7 55.9 73.7 1691/610 81.8 74.9 81.4 57.9 96.8
InternVL2.5-8B 82.6 78.0 80.3 58.9 75.9 1690/610 84.4 76.5 82.7 62.4 97.4
@@
Ablation study on the Segmentation Token Design. We have explored several segmentation token designs, with the results shown in Tab. 9. When using the repeat strategy, which makes the LLM output N [SEG] tokens for N frames, this approach slightly decreases video segmentation performance due to being more prone to missing or repeating [SEG] tokens. We also explored using different segmentation tokens for different frames, such as setting [SEG 1] to [SEG n] to segment the object across frames 1 to n. However, this prevented knowledge sharing with image segmentation tasks, resulting in a larger performance drop.
@@
Table 9: Ablation study on [SEG] token design.
@@ Appendix A More Experiment Results
Appendix B Model Details and Implementation Details
Ref-VOS inference pipeline with SAM-2 knowledge.
@@ 17       Update Memory: $Mem$leftarrow Cross-Attention($\{Mem,M_{i}\}$);
As described in Algorithm 1, the pipeline begins by extracting the first five frames of the input video as keyframes, ensuring that they capture the critical context for the following processing. These key frames are fed into CLIP and flattened to visual sequential tokens for LLM processing. The LLM takes the visual and language tokens as input and uses these tokens to extract information about the video to generate the “[SEG]” token. In SAM-2, the prompt encoder encodes boxes or clicks to prompt embeddings for object referring. Different from SAM-2, we use two linear layers to project the “[SEG]” token into the language prompt embedding, which serves as an extension of the SAM-2 prompt encoders. With the language prompt embedding, we use the SAM-2 decoder to generate the masks of the key frames. Then, we use the memory encoder of SAM-2 to generate a memory based on the output key-frame masks. Finally, the memory attention in SAM-2 generates the remaining masks using the memory generated from the key-frame and previous non-key-frame masks.
Appendix C More Details on Ref-SAV Dataset
@@
Comparison with existing Ref-VOS benchmarks. In Tab. 16, we further compare the previous existing Ref-VOS benchmarks from five different aspects: short text, long text, large object motion, large camera motion, and heavy occlusion.
The previous benchmarks only contain partial aspects of these five challenging cases. On the other hand, our benchmark is built from SAM-2 [67] with long text, which is more challenging than previous benchmarks. This is why the previous Ref-VOS models cannot achieve good results on our benchmark.

### v2

@@ 2 Related Work
Referring Segmentation. This task aims to output specific masks (for image inputs) or tracked masks (for video input) driven by language description. Earlier works [91, 52, 38, 13, 80, 102] explore various fusion modules to achieve better performance. Then, several stronger models [84] adopt transformer-based methods to achieve unified segmentation and tracking in video. Equipped with LLMs, several recent works [40, 99, 83, 65] propose more complex referring tasks, including reasoning referring or joint mask and caption generation. For example, LISA [40] involves reasoning-based segmentation. Then, GLaMM [66] annotates a new dataset and proposes region-level caption and segmentation tasks. Meanwhile, several recent works explore joint instruction tuning on referring segmentation and conversation [85, 101, 8]. Our method extends these studies into the video domain with the help of SAM-2 [67], while keeping stronger performance on both image/video referring tasks and VQA tasks.
Video Segmentation and Grounding. Current video segmentation methods [47, 34, 106] focus on segmenting and tracking pixels in closed sets. Several works [25, 104] explore open-vocabulary settings. However, the concepts are still limited compared with the knowledge space of LLMs. For video grounding, recent works [28] explore LLM into the joint understanding of video and audio. Recent work VISA [85] also explores reasoning video object segmentation. However, it is trained solely on video segmentation datasets and lacks end-to-end training, which limits its capability in other tasks and the potential of scaling up. Our model, Sa2VA, supports fine-grained spatial-temporal modeling of both static (image) and dynamic (video) visual contents and achieves strong performance on various tasks.
@@ 3 Method
Figure 2: Our proposed Sa2VA model. The model first encodes the input texts, visual prompts, images, and videos into token embeddings. These tokens are then processed through a large language model (LLM). The output text tokens are used to generate the “[SEG]” token and associated language outputs. The SAM-2 decoder receives the image and video features from the SAM-2 encoder, along with the “[SEG]” token, to generate corresponding image and video masks.
@@ 3.1 Unifying Multi-task Representations
Image/Video Chat and Grounded Caption Generation. For image and video chat tasks, given input text tokens $T_{i}$ and corresponding images $I_{i}$ or video $V_{i}$, the model outputs answer text $T_{o}$. For grounded caption generation tasks, the model jointly outputs the masks $M_{o}$ and aligned text $T_{o}$.
@@
Unified Task Representation. Existing works solve the above tasks via specific design models or partial unified models (different tasks have different model weights). In this work, we point out that all the above tasks can be unified as a one-shot instruction-tuning process, since we can leverage the LLMs to handle various visual tokens. The overall process can be formulated as the following:
@@
For chat-only tasks, the model only outputs text tokens $T_{o}$. For referring segmentation tasks, the model outputs masks (image) or masklets (video) $M_{o}$. For grounded caption generation tasks, the model outputs both texts and masks. For different tasks, the input visual tokens are changing according to specific tasks. Since $M_{o}$ (whether image masks or masklets) can be controlled with SEG tokens, one SEG token represents one prompt input of SAM-2’s decoder. Thus, it is natural to put all these tasks into one shared encoder-and-decoder framework and perform end-to-end training with one LLM and the SAM decoder.
Figure 3: Data annotation pipeline. Our proposed automatic data annotation pipeline consists of three stages: object-level, scene-level, and video-level text expression annotation. We use different colors in the final expression to highlight the information derived from each stage. Best view on screen and zoom in.
@@ 3.2 Sa2VA Framework
Pre-trained MLLMs. We adopt pre-trained LLaVA-like models as the MLLMs. It contains one visual encoder, one visual projection layer, and one LLM. The visual encoder takes input images, video, and sub-images as inputs. The visual projection layer maps inputs into visual tokens. These tokens, combined with the input text tokens, are the input of LLMs and the LLMs generate the text token prediction based on them. Note that we adopt pre-trained MLLMs following previous works [99, 66, 40, 85] to leverage their strong capability. For both image and video chat datasets, we follow the same pipeline [9, 2] without modification.
@@
Tuning SAM-2 Decoder via SEG Tokens. We connect SAM-2 and the MLLM via the special token “[SEG]”. The hidden states of the “[SEG]” token are used as a new type of prompt and fed into SAM-2’s Decoder, where they are decoded into segmentation masks. The hidden states of “[SEG]” can be seen as a novel spatial-temporal prompt for SAM-2. SAM-2 segments the corresponding object mask in image and video based on the spatial-temporal prompt. During training, the SAM-2 decoder can be tuned to understand the spatial-temporal prompt, and gradients can be backpropagated through the “[SEG]” token to the MLLM, allowing it to output the spatial-temporal prompt better.
Utilizing SAM-2 knowledge for mask tracking. For Ref-VOS tasks, we design a simple framework to achieve strong results on public benchmarks. In particular, for giving input video, we adopt a “[SEG]” token to generate the masks of the key frames. Then, we use the memory encoded by the key frame features to generate the mask for the remaining frames. We detailed this process in the appendix.
@@ 3.3 Ref-SAV Dataset and Benchmark
    Object-level annotation. We first select the frame with the largest object area from the video and mask out the non-object pixels. The cropped and full images are then fed separately into InternVL2-76B [9] to generate detailed descriptions. The descriptions are processed by Qwen2-72B [86] for consistency checking, with conflicting descriptions discarded.
@@
    Scene-level annotation. Yellow contours are used to highlight the object in the image. Both the image and the object-level description from the previous stage are sent into InternVL2-76B [9] to generate a detailed object description that includes relationships to the scene and surrounding objects.
@@
Table 2: Experiment results on image/video referring segmentation benchmarks and image/video chat benchmarks. For MME dataset, A / B means the perception (A) and cognition (B) scores, while scores without “/” mean the total score (A + B)
@@
RefCOCO [36] RefCOCO+ [36] RefCOCOg [90] MeViS [13] Ref-DAVIS17 [38] Ref-YTVOS [69] ReVOS [85] MME [19] MMBench [59] SEED-Bench [41] Video-MME [20] MMBench-Video [17] GCG [66]
LLAVA-1.5-13B [54] - - - - - - - 1531 68.8 70.1 - - -
Video-LLaVA-7B [51] - - - - - - - - 60.9 - 39.9 1.03 -
LLaMA-VID-7B [50] - - - - - - - 1521 65.1 59.9 - 1.08 -
mPLUG-Owl3-8B [89] - - - - - - - - 77.6 - 53.5 1.35 -
InternVL2-8B [9] - - - - - - - - 81.7 76.2 54.0 1.28 -
PixelLM-7B [68] 73.0 66.3 69.3 - - - - 309/135 17.4 - - - -
LaSagnA [77] 76.8 66.4 70.6 - - - - 0/0 0.0 - - - -
LISA-7B [40] 74.1 62.4 66.4 - - - - 1/1 0.4 - - - -
GLaMM-7B [66] 79.5 72.6 74.2 - - - - 14/9 36.8 - - - 28.9
LLaVA-G-7B [96] 77.1 68.8 71.5 - - - - - - - - - -
GSVA-13B [83] 79.2 70.3 75.7 - - - - - - - - - -
OMG-LLaVA-7B [99] 78.0 69.1 72.9 - - - - 1177/235 47.9 56.5 - - 29.9
VISA-13B [85] 72.4 59.8 65.5 44.5 70.4 63.0 50.9 - - - - - -
Sa2VA-1B (Ours) 77.4 69.9 72.3 41.7 72.3 65.3 47.6 1381/405 68.3 64.8 39.9 1.07 23.8
Sa2VA-4B (Ours) 78.9 71.7 74.1 46.2 73.8 70.0 53.2 1536/530 77.3 73.3 50.4 1.23 28.2
Sa2VA-8B (Ours) 81.6 76.2 78.7 46.9 75.2 70.7 57.6 1617/511 81.6 75.1 52.1 1.34 31.0
Sa2VA-26B (Ours) 82.5 78.8 79.7 46.2 77.0 70.1 58.4 1691/538 83.7 76.8 52.6 1.45 33.5
@@ 4 Experiments
Baseline. We construct the baseline by combining the SOTA MLLM models like InternVL2 [9] and SAM2 [67]. Similar to previous works  [40, 66, 99], the segmentation mask is obtained by decoding the hidden state of the “[SEG]” token through SAM2’s decoder. Inspired by Mask2Former-VIS [10], an object shares the same “[SEG]” token throughout the video frames, allowing our model to handle both image and video referring segmentation tasks in a unified way. In addition, we further use more advanced MLLM models, InternVL2-5 to show the model scaling effect of Sa2VA.
@@
Table 4: Performance on image-level benchmarks with MLLMs with segmentation capability. Our method achieves the best accuracy trade-off between image chat and referring segmentation datasets.
@@ 4.2 Ablation Studies
MLLM RefCOCO [36] RefCOCO+ [36] RefCOCOg [90] MeViS (val_u) [13] MeViS (val) [13] Ref-YTVOS [69] Ref-DAVIS17 [38] MME [19] MMBench [59] SEED-Bench [41] AI2D [37] MMStar [7] SQAtest [60]
InternVL2.0-4B 80.4 74.3 76.7 52.1 46.2 70.0 73.8 1553/540 76.8 72.6 79.9 53.7 95.8
InternVL2.0-8B 81.9 76.5 78.9 57.0 46.9 70.7 75.2 1651/578 82.4 75.5 82.1 60.3 96.8
InternVL2.5-1B 79.6 73.6 77.7 53.4 47.0 68.0 69.5 1504/434 71.9 71.0 69.2 48.6 89.9
InternVL2.5-4B 82.4 77.6 79.7 55.9 46.4 71.3 73.7 1691/610 81.8 74.9 81.4 57.9 96.8
InternVL2.5-8B 82.6 78.0 80.3 58.9 51.5 72.3 75.9 1690/610 84.4 76.5 82.7 62.4 97.4
InternVL2.5-26B 82.9 79.3 81.2 61.8 52.1 75.1 78.6 1698/653 85.8 78.3 85.7 67.0 98.4
@@
Ablation study on the Segmentation Token Design. We have explored several segmentation token designs, with the results shown in Tab. 9. When using the repeat strategy, which makes the LLM output N “[SEG]” tokens for N frames, this approach slightly decreases video segmentation performance due to being more prone to missing or repeating “[SEG]” tokens. We also explored using different segmentation tokens for different frames, such as setting “[SEG 1]” to “[SEG n]” to segment the object across frames 1 to n. However, this prevented knowledge sharing with image segmentation tasks, resulting in a larger performance drop.
@@
Table 9: Ablation study on “[SEG]” token design.
@@ 17       Update Memory: $Mem$leftarrow Cross-Attention($\{Mem,M_{i}\}$);
Appendix B Model Details and Implementation Details
@@
Table 17: Statistics comparison with other Ref-VOS datasets.
Datasets Video Object Expression Mask Avg. Len.
Ref-YTVOS 3,978 7,451 15,009 131k 9.68
MeViS 2,006 8,171 28,570 443k 7.07
ReVOS 1,042 9,084 35,074 273k 10.48
Ref-SAV 37,311 72,509 72,509 6.0m 83.6
Ref-VOS inference pipeline with SAM-2 knowledge. As described in Algorithm 1, the pipeline begins by extracting the first five frames of the input video as keyframes, ensuring that they capture the critical context for the following processing. These key frames are fed into CLIP and flattened to visual sequential tokens for LLM processing. The LLM takes the visual and language tokens as input and uses these tokens to extract information about the video to generate the “[SEG]” token. In SAM-2, the prompt encoder encodes boxes or clicks to prompt embeddings for object referring. Different from SAM-2, we use two linear layers to project the “[SEG]” token into the language prompt embedding, which serves as an extension of the SAM-2 prompt encoders. With the language prompt embedding, we use the SAM-2 decoder to generate the masks of the key frames. Then, we use the memory encoder of SAM-2 to generate a memory based on the output key-frame masks. Finally, the memory attention in SAM-2 generates the remaining masks using the memory generated from the key-frame and previous non-key-frame masks.
Appendix C More Details on Ref-SAV Dataset
Comparison with existing Ref-VOS benchmarks. In Tab. 16, we further compare the previous existing Ref-VOS benchmarks from five different aspects: short text, long text, large object motion, large camera motion, and heavy occlusion. As in Table 17, our dataset has more detailed captions (Avg. Len. refers to the average length of the referring description) and many more masks (about 10x). The proposed data can lead to better instruction-following capability for complex descriptions. The previous benchmarks only contain partial aspects of these five challenging cases. On the other hand, our benchmark is built from SAM-2 [67] with long text, which is more challenging than previous benchmarks. This is why the previous Ref-VOS models cannot achieve good results on our benchmark.